# Higher Lie characters and root enumeration in classical Weyl groups

### v1

@@ 6. Root enumerators of type $D$
In this section we prove Theorems 1.2, 1.6 and 1.7. In Subsection 6.1 we define signed root enumerators of type $B$ and state the main result of this section, Theorem 6.6, which is a signed analogue of Theorem 5.3. Its proof uses Theorem 6.7, which provides a generating function for signed root enumerators in $B_{n}$. Theorem 6.7 is stated and proved in Subsection 6.2, and is used in Subsection 6.3, together with Theorem 4.3, to prove Theorem 6.6. Then, in Subsection 6.4, we deduce Theorems 1.6 and 1.2. Finally, in Subsection 6.5, we use Theorem 1.6 to prove Theorem 1.7.
@@ 6.1. Signed root enumerators of type $B$
$\varphi^{B_{n}}(y):=|\{x\in B_{n}\,:\,x^{k}=y\}|=\sum_{\begin{subarray}{c}x\in B_{n}\\ x^{k}=y\end{subarray}}1\qquad(y\in B_{n}).$
@@
${\overline{\varphi}}_{k}^{B_{n}}(y):=\sum_{\begin{subarray}{c}x\in B_{n}\\ x^{k}=y\end{subarray}}\chi(x)\qquad(y\in B_{n}).$
We shall also use the corresponding signed $k$-th root enumerator in ${\mathbb{Z}}_{2}=\{+1,-1\}$,
${\overline{\varphi}}_{k}^{{\mathbb{Z}}_{2}}(\theta):=\sum_{\begin{subarray}{c}\varepsilon\in{\mathbb{Z}}_{2}\\ \varepsilon^{k}=\theta\end{subarray}}\chi(\varepsilon)=\sum_{\begin{subarray}{c}\varepsilon\in{\mathbb{Z}}_{2}\\ \varepsilon^{k}=\theta\end{subarray}}\varepsilon\qquad(\theta\in{\mathbb{Z}}_{2}).$
@@
${\overline{\psi}}_{k}^{B_{n}}:=\sum_{{\underline{\lambda}}\vdash_{k,-}n}\psi^{\underline{\lambda}}.$
@@ 6.2. Generating function for signed root enumerators
In this subsection we state and prove a generating function (Theorem 6.7) for the values of the signed $k$-th root enumerator ${\overline{\varphi}}_{k}^{B_{n}}$ on arbitrary elements of $B_{n}$, in terms of signed $h$-th root enumerators ${\overline{\varphi}}_{h}^{{\mathbb{Z}}_{2}}$ for various divisors $h$ of $k$.
@@
$\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}\frac{{\overline{\varphi}}_{k}^{B_{n}}({\underline{\mu}})}{|Z_{\underline{\mu}}|}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}=\exp\left(\sum_{j,\theta}\sum_{\begin{subarray}{c}h|k\\ \gcd(h,j)=1\end{subarray}}\frac{t_{j,\theta}^{jk/h}}{2jk/h}\cdot{\overline{\varphi}}_{h}^{{\mathbb{Z}}_{2}}(\theta)\right).$
@@
so that $|R_{h,\theta}|=\varphi^{{\mathbb{Z}}_{2}}_{h}(\theta)$.
@@
$\sum_{{\underline{\lambda}}\vdash n}\sum_{\begin{subarray}{c}x\in C_{\underline{\lambda}}\\ x^{k}=y\end{subarray}}\chi(x)={\overline{\varphi}}_{k}^{B_{n}}(y).$
Recall that $\sum_{\varepsilon\in R_{h,\theta}}\varepsilon={\overline{\varphi}}_{h}^{{\mathbb{Z}}_{2}}(\theta)$ and $|Z_{\underline{\mu}}|=|B_{n}|/|C_{\underline{\mu}}|$. Denoting ${\underline{t}}^{{\underline{c}}({\underline{\mu}})}:=\prod_{j,\theta}t_{j,\theta}^{jb_{j,\theta}}$, Theorem 6.9 reduces to Theorem 6.7, presenting the (unrefined) generating function for the signed $k$-th root enumerator in $B_{n}$. ∎
@@ 6.3. Signed root enumerators and higher Lie characters
$\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}\frac{{\overline{\varphi}}_{k}^{B_{n}}({\underline{\mu}})}{|Z_{\underline{\mu}}|}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}=\exp\left(\sum_{j,\theta}\sum_{\begin{subarray}{c}h|k\\ \gcd(h,j)=1\end{subarray}}\frac{t_{j,\theta}^{jk/h}}{2jk/h}\cdot{\overline{\varphi}}_{h}^{{\mathbb{Z}}_{2}}(\theta)\right).$
@@
$\sum_{n\geq 0}\sum_{{\underline{\lambda}}\vdash n}\sum_{{\underline{\mu}}\vdash n}\frac{\psi^{\underline{\lambda}}({\underline{\mu}})}{|Z_{\underline{\mu}}|}{\underline{s}}^{{\underline{c}}({\underline{\lambda}})}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}=\exp\left(\sum_{i,\varepsilon}\sum_{j,\theta}\sum_{e|\gcd(i,j)}\frac{K_{\varepsilon,\theta}(e)(s_{i,\varepsilon}t_{j,\theta})^{ij/e}}{2ij/e}\right),$
@@
$\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}\frac{{\overline{\psi}}_{k}^{B_{n}}({\underline{\mu}})}{|Z_{\underline{\mu}}|}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}$ $=\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}\sum_{{\underline{\lambda}}\vdash_{k,-}n}\frac{\psi^{\underline{\lambda}}({\underline{\mu}})}{|Z_{\underline{\mu}}|}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}$
$=\exp\left(\sum_{j,\theta}\sum_{\begin{subarray}{c}i\\ i|k\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{k/i}=-1\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|\gcd(i,j)\end{subarray}}\frac{K_{\varepsilon,\theta}(e)t_{j,^{\theta}}^{ij/e}}{2ij/e}\right).$
@@
${\overline{\varphi}}_{k}^{B_{n}}={\overline{\psi}}_{k}^{B_{n}}\qquad(\forall n\geq 0)$
@@
(5) $\sum_{\begin{subarray}{c}h\\ h|k\\ \gcd(h,j)=1\end{subarray}}\frac{t_{j,\theta}^{jk/h}}{2jk/h}\cdot{\overline{\varphi}}_{h}^{{\mathbb{Z}}_{2}}(\theta)=\sum_{\begin{subarray}{c}i\\ i|k\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{k/i}=-1\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|\gcd(i,j)\end{subarray}}\frac{K_{\varepsilon,\theta}(e)t_{j,\theta}^{ij/e}}{2ij/e}.$
Letting $d:=i/e$ and then $h:=k/d$ in the RHS of (5) yields
RHS $=\sum_{\begin{subarray}{c}i\\ i|k\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|i\\ e|j\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{k/i}=-1\end{subarray}}\frac{K_{\varepsilon,\theta}(e)t_{j,\theta}^{ij/e}}{2ij/e}$
$=\sum_{\begin{subarray}{c}d\\ d|k\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|(k/d)\\ e|j\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{k/(de)}=-1\end{subarray}}\frac{K_{\varepsilon,\theta}(e)t_{j,\theta}^{jd}}{2jd}$
$=\sum_{\begin{subarray}{c}h\\ h|k\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|h\\ e|j\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{h/e}=-1\end{subarray}}\frac{K_{\varepsilon,\theta}(e)t_{j,\theta}^{jk/h}}{2jk/h}.$
@@
(6) $\sum_{\begin{subarray}{c}e\\ e|\gcd(h,j)\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{h/e}=-1\end{subarray}}K_{\varepsilon,\theta}(e)=\begin{cases}{\overline{\varphi}}_{h}^{{\mathbb{Z}}_{2}}(\theta),&\text{if }\gcd(h,j)=1;\\ 0,&\text{otherwise.}\end{cases}$
@@
If $\gcd(h,j)\neq 1$ then the RHS here is zero, and if $\gcd(h,j)=1$ then the RHS is ${\overline{\varphi}}_{h}^{{\mathbb{Z}}_{2}}(\theta)$. This proves (6) and completes the proof of the theorem. ∎
@@
Fix an integer $k$, a positive integer $n$ and an element $y\in B_{n}$. By the definitions of the root enumerator $\varphi_{k}^{B_{n}}$, the signed root enumerator ${\overline{\varphi}}_{k}^{B_{n}}$ (see Definition 6.3) and the character $\chi$,
$\varphi_{k}^{B_{n}}(y)+{\overline{\varphi}}_{k}^{B_{n}}(y)=\sum_{\begin{subarray}{c}x\in B_{n}\\ x^{k}=y\end{subarray}}1+\sum_{\begin{subarray}{c}x\in B_{n}\\ x^{k}=y\end{subarray}}\chi(x)=\sum_{\begin{subarray}{c}x\in D_{n}\\ x^{k}=y\end{subarray}}2=\begin{cases}2\varphi_{k}^{D_{n}}(y),&\text{if }y\in D_{n};\\ 0,&\text{if }y\in B_{n}\setminus D_{n}.\end{cases}$
@@
$\psi_{k}^{B_{n}}(y)+{\overline{\psi}}_{k}^{B_{n}}(y)=\begin{cases}2\varphi_{k}^{D_{n}}(y),&\text{if }y\in D_{n};\\ 0,&\text{if }y\in B_{n}\setminus D_{n}.\end{cases}$
@@
Consider the expansion of the class function $\varphi_{k}^{D_{n}}$ as a linear combination of irreducible $D_{n}$-characters. By Theorem 1.6, $2\varphi_{k}^{D_{n}}$ is a restriction (to $D_{n}$) of a proper character of $B_{n}$, and is therefore a proper character of $D_{n}$. In particular, all the coefficients in this expansion are nonnegative. By the result of Frobenius cited at the beginning of Section 1, all the coefficients in the expansion of $\varphi_{k}^{D_{n}}$ are integers. This completes the proof. ∎
@@ 6.5. Higher Lie Characters of type $D$
If $H$ is an index $2$ subgroup of a finite group $G$ then, for every odd $k$, $\varphi^{H}_{k}=\varphi^{G}_{k}\downarrow_{H}^{G}$.
@@
Thus, all $k$-th roots (in $G$) of an element of $H$ belong to $H$, namely $\varphi^{H}_{k}=\varphi^{G}_{k}\downarrow_{H}^{G}$. ∎
@@
$\varphi^{D_{n}}_{k}=\varphi^{B_{n}}_{k}\downarrow^{B_{n}}_{D_{n}}=\sum_{C\in\operatorname{Conj}(B_{n})\,:\,C^{k}=\{1\}}\psi^{C}_{B_{n}}\downarrow^{B_{n}}_{D_{n}}.$
@@
$\varphi^{D_{n}}_{k}=\sum_{C\in\operatorname{Conj}(D_{n})\,:\,C^{k}=\{1\}}\psi^{C}_{B_{n}}\downarrow^{B_{n}}_{D_{n}}=\sum_{C\in\operatorname{Conj}(D_{n})\,:\,C^{k}=\{1\}}\psi_{D_{n}}^{C},$
@@
$2\varphi^{D_{n}}_{k}=\sum_{C\in\operatorname{Conj}(B_{n})\,:\,C^{k}=\{1\}}\psi^{C}_{B_{n}}\downarrow_{D_{n}}^{B_{n}}.$
@@
$2\varphi^{D_{n}}_{k}=\sum_{C\in\operatorname{Conj}(D_{n})\,:\,C^{k}=\{1\}}\psi^{C}_{B_{n}}\downarrow_{D_{n}}^{B_{n}}+\sum_{C\in\operatorname{Conj}(D_{n})\,:\,C^{k}=\{1\}}\psi^{w_{0}C}_{B_{n}}\downarrow_{D_{n}}^{B_{n}}.$
@@
$2\varphi^{D_{n}}_{k}=\sum_{C\in\operatorname{Conj}(D_{n})\,:\,C^{k}=\{1\}}(\psi^{C}_{B_{n}}(1+\chi))\downarrow_{D_{n}}^{B_{n}}=\sum_{C\in\operatorname{Conj}(D_{n})\,:\,C^{k}=\{1\}}2\psi^{C}_{B_{n}}\downarrow_{D_{n}}^{B_{n}},$

### v2

@@ 6. Root enumerators of type $D$
In this section we prove Theorems 1.2, 1.6 and 1.8. In Subsection 6.1 we define signed root enumerators of type $B$ and state the main result of this section, Theorem 6.6, which is a signed analogue of Theorem 5.3. Its proof uses Theorem 6.7, which provides a generating function for signed root enumerators in $B_{n}$. Theorem 6.7 is stated and proved in Subsection 6.2, and is used in Subsection 6.3, together with Theorem 4.3, to prove Theorem 6.6. Then, in Subsection 6.4, we deduce Theorems 1.6 and 1.2. Finally, in Subsection 6.5, we use Theorem 1.6 to prove Theorem 1.8.
@@ 6.1. Signed root enumerators of type $B$
$\rho_{k}^{B_{n}}(y):=|\{x\in B_{n}\,:\,x^{k}=y\}|=\sum_{\begin{subarray}{c}x\in B_{n}\\ x^{k}=y\end{subarray}}1\qquad(y\in B_{n}).$
@@
${\overline{\rho}}_{k}^{B_{n}}(y):=\sum_{\begin{subarray}{c}x\in B_{n}\\ x^{k}=y\end{subarray}}\chi(x)\qquad(y\in B_{n}).$
We shall also use the signed $k$-th root enumerator in ${\mathbb{Z}}_{2}=\{+1,-1\}$,
${\overline{\rho}}_{k}^{{\mathbb{Z}}_{2}}(\theta):=\sum_{\begin{subarray}{c}\varepsilon\in{\mathbb{Z}}_{2}\\ \varepsilon^{k}=\theta\end{subarray}}\chi(\varepsilon)=\sum_{\begin{subarray}{c}\varepsilon\in{\mathbb{Z}}_{2}\\ \varepsilon^{k}=\theta\end{subarray}}\varepsilon\qquad(\theta\in{\mathbb{Z}}_{2}).$
@@
${\overline{\psi}}_{k}^{B_{n}}:=\sum_{{\underline{\lambda}}\vdash_{k,-}n}\psi_{B_{n}}^{\underline{\lambda}}.$
@@ 6.2. Generating function for signed root enumerators
In this subsection we state and prove a generating function (Theorem 6.7) for the values of the signed $k$-th root enumerator ${\overline{\rho}}_{k}^{B_{n}}$ on arbitrary elements of $B_{n}$, in terms of signed $h$-th root enumerators ${\overline{\rho}}_{h}^{{\mathbb{Z}}_{2}}$ for various divisors $h$ of $k$.
@@
$\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}{\overline{\rho}}_{k}^{B_{n}}({\underline{\mu}})\,\frac{{\underline{t}}^{{\underline{c}}({\underline{\mu}})}}{|Z_{\underline{\mu}}|}=\exp\left(\sum_{j,\theta}\sum_{\begin{subarray}{c}h|k\\ \gcd(h,j)=1\end{subarray}}{\overline{\rho}}_{h}^{{\mathbb{Z}}_{2}}(\theta)\,\frac{t_{j,\theta}^{jk/h}}{2jk/h}\right).$
@@
so that $|R_{h,\theta}|=\rho^{{\mathbb{Z}}_{2}}_{h}(\theta)$.
@@
$\sum_{{\underline{\lambda}}\vdash n}\sum_{\begin{subarray}{c}x\in C_{\underline{\lambda}}\\ x^{k}=y\end{subarray}}\chi(x)={\overline{\rho}}_{k}^{B_{n}}(y).$
Recall that $\sum_{\varepsilon\in R_{h,\theta}}\varepsilon={\overline{\rho}}_{h}^{{\mathbb{Z}}_{2}}(\theta)$ and $|Z_{\underline{\mu}}|=|B_{n}|/|C_{\underline{\mu}}|$. Denoting ${\underline{t}}^{{\underline{c}}({\underline{\mu}})}:=\prod_{j,\theta}t_{j,\theta}^{jb_{j,\theta}}$, Theorem 6.9 reduces to Theorem 6.7, presenting the (unrefined) generating function for the signed $k$-th root enumerator in $B_{n}$. ∎
@@ 6.3. Signed root enumerators and higher Lie characters
$\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}{\overline{\rho}}_{k}^{B_{n}}({\underline{\mu}})\,\frac{{\underline{t}}^{{\underline{c}}({\underline{\mu}})}}{|Z_{\underline{\mu}}|}=\exp\left(\sum_{j,\theta}\sum_{\begin{subarray}{c}h|k\\ \gcd(h,j)=1\end{subarray}}{\overline{\rho}}_{h}^{{\mathbb{Z}}_{2}}(\theta)\,\frac{t_{j,\theta}^{jk/h}}{2jk/h}\right).$
@@
$\sum_{n\geq 0}\sum_{{\underline{\lambda}}\vdash n}\sum_{{\underline{\mu}}\vdash n}\psi_{B_{n}}^{\underline{\lambda}}({\underline{\mu}})\,\frac{{\underline{s}}^{{\underline{c}}({\underline{\lambda}})}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}}{|Z_{\underline{\mu}}|}=\exp\left(\sum_{i,\varepsilon}\sum_{j,\theta}\sum_{e|\gcd(i,j)}K_{\varepsilon,\theta}(e)\,\frac{(s_{i,\varepsilon}t_{j,\theta})^{ij/e}}{2ij/e}\right),$
@@
$\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}{\overline{\psi}}_{k}^{B_{n}}({\underline{\mu}})\,\frac{{\underline{t}}^{{\underline{c}}({\underline{\mu}})}}{|Z_{\underline{\mu}}|}$ $=\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}\sum_{{\underline{\lambda}}\vdash_{k,-}n}\psi_{B_{n}}^{\underline{\lambda}}({\underline{\mu}})\,\frac{{\underline{t}}^{{\underline{c}}({\underline{\mu}})}}{|Z_{\underline{\mu}}|}$
$=\exp\left(\sum_{j,\theta}\sum_{\begin{subarray}{c}i\\ i|k\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{k/i}=-1\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|\gcd(i,j)\end{subarray}}K_{\varepsilon,\theta}(e)\,\frac{t_{j,^{\theta}}^{ij/e}}{2ij/e}\right).$
@@
${\overline{\rho}}_{k}^{B_{n}}={\overline{\psi}}_{k}^{B_{n}}\qquad(\forall n\geq 0)$
@@
(5) $\sum_{\begin{subarray}{c}h\\ h|k\\ \gcd(h,j)=1\end{subarray}}{\overline{\rho}}_{h}^{{\mathbb{Z}}_{2}}(\theta)\,\frac{t_{j,\theta}^{jk/h}}{2jk/h}=\sum_{\begin{subarray}{c}i\\ i|k\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{k/i}=-1\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|\gcd(i,j)\end{subarray}}K_{\varepsilon,\theta}(e)\,\frac{t_{j,\theta}^{ij/e}}{2ij/e}.$
Letting $d:=i/e$ and then $h:=k/d$ on the RHS of (5) yields
RHS $=\sum_{\begin{subarray}{c}i\\ i|k\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|i\\ e|j\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{k/i}=-1\end{subarray}}K_{\varepsilon,\theta}(e)\,\frac{t_{j,\theta}^{ij/e}}{2ij/e}$
$=\sum_{\begin{subarray}{c}d\\ d|k\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|(k/d)\\ e|j\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{k/(de)}=-1\end{subarray}}K_{\varepsilon,\theta}(e)\,\frac{t_{j,\theta}^{jd}}{2jd}$
$=\sum_{\begin{subarray}{c}h\\ h|k\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|h\\ e|j\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{h/e}=-1\end{subarray}}K_{\varepsilon,\theta}(e)\,\frac{t_{j,\theta}^{jk/h}}{2jk/h}.$
@@
(6) $\sum_{\begin{subarray}{c}e\\ e|\gcd(h,j)\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{h/e}=-1\end{subarray}}K_{\varepsilon,\theta}(e)=\begin{cases}{\overline{\rho}}_{h}^{{\mathbb{Z}}_{2}}(\theta),&\text{if }\gcd(h,j)=1;\\ 0,&\text{otherwise.}\end{cases}$
@@
If $\gcd(h,j)\neq 1$ then the RHS here is zero, and if $\gcd(h,j)=1$ then the RHS is ${\overline{\rho}}_{h}^{{\mathbb{Z}}_{2}}(\theta)$. This proves (6) and completes the proof of the theorem. ∎
@@
Fix an integer $k$, a positive integer $n$ and an element $y\in B_{n}$. By the definitions of the root enumerator $\rho_{k}^{B_{n}}$, the signed root enumerator ${\overline{\rho}}_{k}^{B_{n}}$ (see Definition 6.3) and the character $\chi$,
$\rho_{k}^{B_{n}}(y)+{\overline{\rho}}_{k}^{B_{n}}(y)=\sum_{\begin{subarray}{c}x\in B_{n}\\ x^{k}=y\end{subarray}}1+\sum_{\begin{subarray}{c}x\in B_{n}\\ x^{k}=y\end{subarray}}\chi(x)=\sum_{\begin{subarray}{c}x\in D_{n}\\ x^{k}=y\end{subarray}}2=\begin{cases}2\rho_{k}^{D_{n}}(y),&\text{if }y\in D_{n};\\ 0,&\text{if }y\in B_{n}\setminus D_{n}.\end{cases}$
@@
$\psi_{k}^{B_{n}}(y)+{\overline{\psi}}_{k}^{B_{n}}(y)=\begin{cases}2\rho_{k}^{D_{n}}(y),&\text{if }y\in D_{n};\\ 0,&\text{if }y\in B_{n}\setminus D_{n}.\end{cases}$
@@
Consider the expansion of the class function $\rho_{k}^{D_{n}}$ as a linear combination of irreducible $D_{n}$-characters. By Theorem 1.6, $2\rho_{k}^{D_{n}}$ is a restriction (to $D_{n}$) of a proper character of $B_{n}$, and is therefore a proper character of $D_{n}$. In particular, all the coefficients in this expansion are nonnegative. By the result of Frobenius cited at the beginning of Section 1, all the coefficients in the expansion of $\rho_{k}^{D_{n}}$ are integers. This completes the proof. ∎
@@ 6.5. Higher Lie Characters of type $D$
If $H$ is an index $2$ subgroup of a finite group $G$ then, for every odd $k$, $\rho^{H}_{k}=\rho^{G}_{k}\downarrow_{H}^{G}$.
@@
Thus, all $k$-th roots (in $G$) of an element of $H$ belong to $H$, namely $\rho^{H}_{k}=\rho^{G}_{k}\downarrow_{H}^{G}$. ∎
@@
$\rho^{D_{n}}_{k}=\rho^{B_{n}}_{k}\downarrow^{B_{n}}_{D_{n}}=\sum_{C\in\operatorname{Conj}(B_{n})\,:\,C^{k}=\{1\}}\psi_{B_{n}}^{C}\downarrow^{B_{n}}_{D_{n}}.$
@@
$\rho^{D_{n}}_{k}=\sum_{C\in\operatorname{Conj}(D_{n})\,:\,C^{k}=\{1\}}\psi_{B_{n}}^{C}\downarrow^{B_{n}}_{D_{n}}=\sum_{C\in\operatorname{Conj}(D_{n})\,:\,C^{k}=\{1\}}\psi_{D_{n}}^{C},$
@@
$2\rho^{D_{n}}_{k}=\sum_{C\in\operatorname{Conj}(B_{n})\,:\,C^{k}=\{1\}}\psi_{B_{n}}^{C}\downarrow_{D_{n}}^{B_{n}}.$
@@
$2\rho^{D_{n}}_{k}=\sum_{C\in\operatorname{Conj}(D_{n})\,:\,C^{k}=\{1\}}\psi_{B_{n}}^{C}\downarrow_{D_{n}}^{B_{n}}+\sum_{C\in\operatorname{Conj}(D_{n})\,:\,C^{k}=\{1\}}\psi_{B_{n}}^{w_{0}C}\downarrow_{D_{n}}^{B_{n}}.$
@@
$2\rho^{D_{n}}_{k}=\sum_{C\in\operatorname{Conj}(D_{n})\,:\,C^{k}=\{1\}}(\psi^{C}_{B_{n}}(1+\chi))\downarrow_{D_{n}}^{B_{n}}=\sum_{C\in\operatorname{Conj}(D_{n})\,:\,C^{k}=\{1\}}2\psi^{C}_{B_{n}}\downarrow_{D_{n}}^{B_{n}},$